Logiweb bibliography of kvanti

Up Help

[0] kvanti

[1] base

[2] check

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-12-15.UTC:00:19:10.164930 = MJD-54084.TAI:00:19:43.164930 = LGT-4672858783164930e-6